OpenStudy (anonymous):

reduce this algebraic fraction: a^3+a^2b/15a X 25/3b+3a

OpenStudy (aum):

\( \Large \frac{a^3+a^2b}{15a} \times \frac{25}{3b+3a} = \frac{a^2(a+b)}{15a} \times \frac{25}{3(a+b)} \) Simplify.